OpenStudy (anonymous):

Solve. ( I have NO clue ware to even begin on this one...) (problum will be drawn in the comments in just a bit)

OpenStudy (anonymous):



OpenStudy (anonymous):

How would you undo the -1/3?

OpenStudy (anonymous):

I have no clue D: thats why I need help

OpenStudy (anonymous):

@Emmalineh123 @Manny77

OpenStudy (anonymous):

Allow me to rephrase that. In order to undo a fraction multplied by a variable (-1/3b) what would we multiply it by in order to get the -1/3b to be 1b?

OpenStudy (anonymous):

help please? Q~Q

OpenStudy (anonymous):

Are you asking for help or for answers? Because this website is for help, not for answers.

OpenStudy (anonymous):

uuuh for help to get the answer of course o_O

OpenStudy (anonymous):

Then you will have to contribute to it. Think about what you'd need to multiply -1/3 by to equal 1.

OpenStudy (anonymous):

We aren't going to be here when you have a test to take. So you need to be able to do it by yourself. I'm helping you figure it out.

OpenStudy (anonymous):

i think you should start simpler. if you want 1/3 to become 1 how many thirds are in a whole

OpenStudy (anonymous):

3 O:

OpenStudy (anonymous):

right, you need three times as many thirds to make one. so if you want 1 whole B, then how many times more 1/3B's do you need.

OpenStudy (anonymous):

3 O:...?

OpenStudy (anonymous):

exactly. now a rule of algebra is that whatever is done to one side of the equal sign has to be done to the other. an example 1=1, so if you want to add another to the right hand side you need to add 1 to the other as well to keep it true like 1+1 = 1+1 same with multiplying. if you multiply 2 on one side you need to multiply 2 on the other

OpenStudy (anonymous):

2*1 =2*1. with your problem you just said you need 3 times as many B's on the left hand side. but if you triple one side what would happen to the other

OpenStudy (anonymous):

hmmm so according to what your saying if I understand correctly b is 3...?

OpenStudy (anonymous):

not exactly, im saying if you want to solve for B you need to get one whole B on one side.

OpenStudy (anonymous):

and you said there are 3 thirds in a whole. so to get one third into a whole, then you have to multiply by 3. B doesnt equal 3,

OpenStudy (anonymous):

so how can you get -1/3B = 9, to be -B =(this side of the equation)

OpenStudy (anonymous):

the right hand side still needs solved for and we arent there yet, we are just working the left hand side.

OpenStudy (anonymous):

27 O-O... 27???? :OOO ... 27 >:DDDD

OpenStudy (anonymous):

very good. but. one last thing

OpenStudy (anonymous):

now we have a -B = 27. so if negative B is Positive 27, what is Positive B?

OpenStudy (anonymous):

wow wow wow. hold the phone O_____O that dose not make sense to me. why are you saying "-B" is be negative because it is on the same side of the equation as the"-1/3"...

OpenStudy (anonymous):

so we got that 3 one thirds make a whole, but we have -1/3. so to get a whole here we need 3 negative1/3's, and that would equal -1

OpenStudy (anonymous):

and wouldnt that make the positive b -27? o.O

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

the key to any of these problems is to get your variable, in this case here B, to be a whole number so that it is easier to work with. so we had -1/3B, and we needed to get it to B some how. so the first step we took was to multiply it by 3 to make a whole. that got us to -B. and remember earlier i said what you do to one side you do to the other. (3)* -1/3B = 9(3) <---- here we multiplied both sides by 3. that is how you got the 27. and 3*-1/3 = -1 so -1B = 27 now we want it to be positive B, so to get a negative to be positive you multiply it by a negative, isnt that right. (-1)* -1B = 27* (-1) <---- what we do to one side we do to the other, so multiplying both sides by -1 B = -27

OpenStudy (anonymous):

thank you :DD that''s the first time anyone has put forth that much effort to help me out on open study.

OpenStudy (anonymous):

its no trouble. i just hope it helped :)

OpenStudy (anonymous):

it did :'D!